### v1

# 00footnotetext: Received xxxx June xxxx

Abstract
Using 20.3 $\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at a center-of-mass energy of 3.773 $\,\mathrm{GeV}$ with the BESIII detector, we report an improved search for the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$. An upper limit on its partial branching fraction for photon energies $E_{\gamma}>10\leavevmode\nobreak\ \,\mathrm{MeV}$ is determined to be $1.2\times 10^{-5}$ at 90% confidence level, which excludes most current theoretical predictions. A sophisticated deep learning approach with thorough validation, based on the Transformer architecture, is implemented to efficiently distinguish the signal from massive backgrounds.
keywords:
Charmed hadron, Radiative leptonic decay, BESIII experiment, Deep learning
M. Ablikim(麦迪娜)1, M. N. Achasov4,c, P. Adlarson76, X. C. Ai(艾小聪)81, R. Aliberti35, A. Amoroso75A,75C, Q. An(安琪)72,58,a, Y. Bai(白羽)57, O. Bakina36, Y. Ban(班勇)46,h, H.-R. Bao(包浩然)64, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi75A,75C, E. Bianco75A,75C, A. Bortone75A,75C, I. Boyko36, R. A. Briere5, A. Brueggemann69, H. Cai(蔡浩)77, M. H. Cai(蔡铭航)38,k,l, X. Cai(蔡啸)1,58, A. Calcaterra28A, G. F. Cao(曹国富)1,64, N. Cao(曹宁)1,64, S. A. Cetin62A, X. Y. Chai(柴新宇)46,h, J. F. Chang(常劲帆)1,58, G. R. Che(车国荣)43, Y. Z. Che(车逾之)1,58,64, G. Chelkov36,b, C. Chen(陈琛)43, C. H. Chen(陈春卉)9, Chao Chen(陈超)55, G. Chen(陈刚)1, H. S. Chen(陈和生)1,64, H. Y. Chen(陈弘扬)20, M. L. Chen(陈玛丽)1,58,64, S. J. Chen(陈申见)42, S. L. Chen(陈思璐)45, S. M. Chen(陈少敏)61, T. Chen(陈通)1,64, X. R. Chen(陈旭荣)31,64, X. T. Chen(陈肖婷)1,64, X. Y. Chen(陈心宇)12,g, Y. B. Chen(陈元柏)1,58, Y. Q. Chen34, Z. J. Chen(陈卓俊)25,i, Z. K. Chen(陈梓康)59, S. K. Choi10, X.  Chu(初晓)12,g, G. Cibinetto29A, F. Cossio75C, J. J. Cui(崔佳佳)50, H. L. Dai(代洪亮)1,58, J. P. Dai(代建平)79, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng(邓创旗)73, Z. Y. Deng(邓子艳)1, A. Denig35, I. Denysenko36, M. Destefanis75A,75C, F. De Mori75A,75C, B. Ding(丁彪)67,1, X. X. Ding(丁晓萱)46,h, Y. Ding(丁逸)34, Y. Ding(丁勇)40, Y. X. Ding(丁玉鑫)30, J. Dong(董静)1,58, L. Y. Dong(董燎原)1,64, M. Y. Dong(董明义)1,58,64, X. Dong(董翔)77, M. C. Du(杜蒙川)1, S. X. Du(杜少旭)12,g, S. X. Du(杜书先)81, Y. Y. Duan(段尧予)55, Z. H. Duan(段宗欢)42, P. Egorov36,b, G. F. Fan(樊高峰)42, J. J. Fan(樊俊杰)19, Y. H. Fan(范宇晗)45, J. Fang(方进)59, J. Fang(方建)1,58, S. S. Fang(房双世)1,64, W. X. Fang(方文兴)1, Y. Q. Fang(方亚泉)1,58, R. Farinelli29A, L. Fava75B,75C, F. Feldbauer3, G. Felici28A, C. Q. Feng(封常青)72,58, Y. T. Feng(冯玙潼)72,58, M. Fritsch3, C. D. Fu(傅成栋)1, J. L. Fu(傅金林)64, Y. W. Fu(傅亦威)1,64, H. Gao(高涵)64, X. B. Gao(高鑫博)41, Y. Gao(高扬)72,58, Y. N. Gao(高原宁)46,h, Y. N. Gao(高语浓)19, Y. Y. Gao(高洋洋)30, S. Garbolino75C, I. Garzia29A,29B, P. T. Ge(葛潘婷)19, Z. W. Ge(葛振武)42, C. Geng(耿聪)59, E. M. Gersabeck68, A. Gilman70, K. Goetzen13, J. D. Gong(龚家鼎)34, L. Gong(龚丽)40, W. X. Gong(龚文煊)1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco75A,75C, M. H. Gu(顾旻皓)1,58, Y. T. Gu(顾运厅)15, C. Y. Guan(关春懿)1,64, A. Q. Guo(郭爱强)31, L. B. Guo(郭立波)41, M. J. Guo(国梦娇)50, R. P. Guo(郭如盼)49, Y. P. Guo(郭玉萍)12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han(韩坤霖)64, T. T. Han(韩婷婷)1, F. Hanisch3, K. D. Hao(郝科迪)72,58, X. Q. Hao(郝喜庆)19, F. A. Harris66, K. K. He(何凯凯)55, K. L. He(何康林)1,64, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng(衡月昆)1,58,64, C. Herold60, T. Holtmann3, P. C. Hong(洪鹏程)34, G. Y. Hou(侯国一)1,64, X. T. Hou(侯贤涛)1,64, Y. R. Hou(侯颖锐)64, Z. L. Hou(侯治龙)1, H. M. Hu(胡海明)1,64, J. F. Hu(胡继峰)56,j, Q. P. Hu(胡启鹏)72,58, S. L. Hu(胡圣亮)12,g, T. Hu(胡涛)1,58,64, Y. Hu(胡誉)1, Z. M. Hu(胡忠敏)59, G. S. Huang(黄光顺)72,58, K. X. Huang(黄凯旋)59, L. Q. Huang(黄麟钦)31,64, P. Huang(黄盼)42, X. T. Huang(黄性涛)50, Y. P. Huang(黄燕萍)1, Y. S. Huang(黄永盛)59, T. Hussain74, N. Hüsken35, N. in der Wiesche69, J. Jackson27, Q. Ji(纪全)1, Q. P. Ji(姬清平)19, W. Ji(季旺)1,64, X. B. Ji(季晓斌)1,64, X. L. Ji(季筱璐)1,58, Y. Y. Ji(吉钰瑶)50, Z. K. Jia(贾泽坤)72,58, D. Jiang(姜地)1,64, H. B. Jiang(姜候兵)77, P. C. Jiang(蒋沛成)46,h, S. J. Jiang(蒋思婧)9, T. J. Jiang(蒋庭俊)16, X. S. Jiang(江晓山)1,58,64, Y. Jiang(蒋艺)64, J. B. Jiao(焦健斌)50, J. K. Jiao(焦俊坤)34, Z. Jiao(焦铮)23, S. Jin(金山)42, Y. Jin(金毅)67, M. Q. Jing(荆茂强)1,64, X. M. Jing(景新媚)64, T. Johansson76, S. Kabana33, N. Kalantar-Nayestanaki65, X. L. Kang(康晓琳)9, X. S. Kang(康晓珅)40, M. Kavatsyuk65, B. C. Ke(柯百谦)81, V. Khachatryan27, A. Khoukaz69, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui(奎贤)1,64, N.  Kumar26, A. Kupsc44,76, W. Kühn37, Q. Lan(兰强)73, W. N. Lan(兰文宁)19, T. T. Lei(雷天天)72,58, M. Lellmann35, T. Lenz35, C. Li(李聪)43, C. Li(李澄)72,58, C. Li(李翠)47, C. H. Li(李春花)39, C. K. Li(李春凯)20, D. M. Li(李德民)81, F. Li(李飞)1,58, G. Li(李刚)1, H. B. Li(李海波)1,64, H. J. Li(李惠静)19, H. N. Li(李衡讷)56,j, Hui Li(李慧)43, J. R. Li(李嘉荣)61, J. S. Li(李静舒)59, K. Li(李科)1, K. L. Li(李凯璐)38,k,l, K. L. Li(李凯璐)19, L. J. Li(李林健)1,64, Lei Li(李蕾)48, M. H. Li(李明浩)43, M. R. Li(李明润)1,64, P. L. Li(李佩莲)64, P. R. Li(李培荣)38,k,l, Q. M. Li(李启铭)1,64, Q. X. Li(李起鑫)50, R. Li( 李燃)17,31, S. X. Li(李素娴)12, T.  Li(李腾)50, T. Y. Li(李天佑)43, W. D. Li(李卫东)1,64, W. G. Li(李卫国)1,a, X. Li(李旭)1,64, X. H. Li(李旭红)72,58, X. L. Li(李晓玲)50, X. Y. Li(李晓宇)1,8, X. Z. Li(李绪泽)59, Y. Li(李洋)19, Y. G. Li(李彦谷)46,h, Y. P. Li(李雁鹏)34, Z. J. Li(李志军)59, Z. Y. Li(李紫阳)79, C. Liang(梁畅)42, H. Liang(梁昊)72,58, Y. F. Liang(梁勇飞)54, Y. T. Liang(梁羽铁)31,64, G. R. Liao(廖广睿)14, L. B. Liao(廖立波)59, M. H. Liao(廖明华)59, Y. P. Liao(廖一朴)1,64, J. Libby26, A.  Limphirat60, C. C. Lin(蔺长城)55, C. X. Lin(林创新)64, D. X. Lin(林德旭)31,64, L. Q. Lin(邵麟筌)39, T. Lin(林韬)1, B. J. Liu(刘北江)1, B. X. Liu(刘宝鑫)77, C. Liu(刘成)34, C. X. Liu(刘春秀)1, F. Liu(刘芳)1, F. H. Liu(刘福虎)53, Feng Liu(刘峰)6, G. M. Liu(刘国明)56,j, H. Liu(刘昊)38,k,l, H. B. Liu(刘宏邦)15, H. H. Liu(刘欢欢)1, H. M. Liu(刘怀民)1,64, Huihui Liu(刘汇慧)21, J. B. Liu(刘建北)72,58, J. J. Liu(刘佳佳)20, K. Liu(刘凯)38,k,l, K.  Liu(刘坤)73, K. Y. Liu(刘魁勇)40, Ke Liu(刘珂)22, L. Liu(刘亮)72,58, L. C. Liu(刘良辰)43, Lu Liu(刘露)43, M. H. Liu(刘美宏)12,g, P. L. Liu(刘佩莲)1, Q. Liu(刘倩)64, S. B. Liu(刘树彬)72,58, T. Liu(刘桐)12,g, W. K. Liu(刘维克)43, W. M. Liu(刘卫民)72,58, W. T. Liu(刘婉婷)39, X. Liu(刘翔)38,k,l, X. Liu(刘鑫)39, X. L. Liu(刘兴淋)12,g, X. Y. Liu(刘雪吟)77, Y. Liu(刘英)38,k,l, Y. Liu(刘义)81, Y. Liu(刘媛)81, Y. B. Liu(刘玉斌)43, Z. A. Liu(刘振安)1,58,64, Z. D. Liu(刘宗德)9, Z. Q. Liu(刘智青)50, X. C. Lou(娄辛丑)1,58,64, F. X. Lu(卢飞翔)59, H. J. Lu(吕海江)23, J. G. Lu(吕军光)1,58, Y. Lu(卢宇)7, Y. H. Lu(卢泱宏)1,64, Y. P. Lu(卢云鹏)1,58, Z. H. Lu(卢泽辉)1,64, C. L. Luo(罗成林)41, J. R. Luo(罗家瑞)59, J. S. Luo(罗家顺)1,64, M. X. Luo(罗民兴)80, T. Luo(罗涛)12,g, X. L. Luo(罗小兰)1,58, Z. Y. Lv(吕在裕)22, X. R. Lyu(吕晓睿)64,p, Y. F. Lyu(吕翌丰)43, Y. H. Lyu(吕云鹤)81, F. C. Ma(马凤才)40, H. Ma(马衡)79, H. L. Ma(马海龙)1, J. L. Ma(马俊力)1,64, L. L. Ma(马连良)50, L. R. Ma(马立瑞)67, Q. M. Ma(马秋梅)1, R. Q. Ma(马润秋)1,64, R. Y. Ma(马若云)19, T. Ma(马腾)72,58, X. T. Ma(马晓天)1,64, X. Y. Ma(马骁妍)1,58, Y. M. Ma(马玉明)31, F. E. Maas18, I. MacKay70, M. Maggiora75A,75C, S. Malde70, Q. A. Malik74, Y. J. Mao(冒亚军)46,h, Z. P. Mao(毛泽普)1, S. Marcello75A,75C, F. M. Melendi29A,29B, Y. H. Meng(孟琰皓)64, Z. X. Meng(孟召霞)67, J. G. Messchendorp13,65, G. Mezzadri29A, H. Miao(妙晗)1,64, T. J. Min(闵天觉)42, R. E. Mitchell27, X. H. Mo(莫晓虎)1,58,64, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie(聂麟苏)20, I. B. Nikolaev4,c, Z. Ning(宁哲)1,58, S. Nisar11,m, Q. L. Niu(牛祺乐)38,k,l, W. D. Niu(牛文迪)12,g, S. L. Olsen10,64, Q. Ouyang(欧阳群)1,58,64, S. Pacetti28B,28C, X. Pan(潘祥)55, Y. Pan(潘越)57, A. Pathak10, Y. P. Pei(裴宇鹏)72,58, M. Pelizaeus3, H. P. Peng(彭海平)72,58, Y. Y. Peng(彭云翊)38,k,l, K. Peters13,e, J. L. Ping(平加伦)41, R. G. Ping(平荣刚)1,64, S. Plura35, F. Z. Qi(齐法制)1, H. R. Qi(漆红荣)61, M. Qi(祁鸣)42, S. Qian(钱森)1,58, W. B. Qian(钱文斌)64, C. F. Qiao(乔从丰)64, J. H. Qiao(乔佳辉)19, J. J. Qin(秦佳佳)73, J. L. Qin(覃嘉良)55, L. Q. Qin(秦丽清)14, L. Y. Qin(秦龙宇)72,58, P. B. Qin(秦鹏勃)73, X. P. Qin(覃潇平)12,g, X. S. Qin(秦小帅)50, Z. H. Qin(秦中华)1,58, J. F. Qiu(邱进发)1, Z. H. Qu(屈子皓)73, C. F. Redmer35, A. Rivetti75C, M. Rolo75C, G. Rong(荣刚)1,64, S. S. Rong(荣少石)1,64, F. Rosini28B,28C, Ch. Rosner18, M. Q. Ruan(阮曼奇)1,58, S. N. Ruan(阮氏宁)43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning76, M. Scodeggio29A, K. Y. Shan(尚科羽)12,g, W. Shan(单葳)24, X. Y. Shan(单心钰)72,58, Z. J. Shang(尚子杰)38,k,l, J. F. Shangguan(上官剑锋)16, L. G. Shao(邵立港)1,64, M. Shao(邵明)72,58, C. P. Shen(沈成平)12,g, H. F. Shen(沈宏飞)1,8, W. H. Shen(沈文涵)64, X. Y. Shen(沈肖雁)1,64, B. A. Shi(施伯安)64, H. Shi(史华)72,58, J. L. Shi(石家磊)12,g, J. Y. Shi(石京燕)1, S. Y. Shi(史书宇)73, X. Shi(史欣)1,58, H. L. Song(宋海林)72,58, J. J. Song(宋娇娇)19, T. Z. Song(宋天资)59, W. M. Song(宋维民)34, Y.  J. Song(宋宇镜)12,g, Y. X. Song(宋昀轩)46,h,n, S. Sosio75A,75C, S. Spataro75A,75C, F. Stieler35, S. S Su(苏闪闪)40, Y. J. Su(粟杨捷)64, G. B. Sun(孙光豹)77, G. X. Sun(孙功星)1, H. Sun(孙昊)64, H. K. Sun(孙浩凯)1, J. F. Sun(孙俊峰)19, K. Sun(孙开)61, L. Sun(孙亮)77, S. S. Sun(孙胜森)1,64, T. Sun51,f, Y. C. Sun(孙雨长)77, Y. H. Sun(孙益华)30, Y. J. Sun(孙勇杰)72,58, Y. Z. Sun(孙永昭)1, Z. Q. Sun(孙泽群)1,64, Z. T. Sun(孙振田)50, C. J. Tang(唐昌建)54, G. Y. Tang(唐光毅)1, J. Tang(唐健)59, J. J. Tang(唐嘉骏)72,58, L. F. Tang(唐林发)39, Y. A. Tang(唐迎澳)77, L. Y. Tao(陶璐燕)73, M. Tat70, J. X. Teng(滕佳秀)72,58, J. Y. Tian(田济源)72,58, W. H. Tian(田文辉)59, Y. Tian(田野)31, Z. F. Tian(田喆飞)77, I. Uman62B, B. Wang(王斌)1, B. Wang(王博)59, Bo Wang(王博)72,58, C.  Wang(王超)19, Cong Wang(王聪)22, D. Y. Wang(王大勇)46,h, H. J. Wang(王泓鉴)38,k,l, J. J. Wang(王家驹)77, K. Wang(王科)1,58, L. L. Wang(王亮亮)1, L. W. Wang(王璐仪)34, M. Wang(王萌)50, M.  Wang72,58, N. Y. Wang(王南洋)64, S. Wang(王顺)12,g, T.  Wang(王婷)12,g, T. J. Wang(王腾蛟)43, W.  Wang(王维)73, W. Wang(王为)59, W. P. Wang(王维平)35,58,72,o, X. Wang(王轩)46,h, X. F. Wang(王雄飞)38,k,l, X. J. Wang(王希俊)39, X. L. Wang(王小龙)12,g, X. N. Wang(王新南)1, Y. Wang(王亦)61, Y. D. Wang(王雅迪)45, Y. F. Wang(王贻芳)1,8,64, Y. H. Wang(王英豪)38,k,l, Y. J. Wang(王祎景)72,58, Y. L. Wang(王艺龙)19, Y. N. Wang(王燕宁)77, Y. Q. Wang(王雨晴)1, Yaqian Wang(王亚乾)17, Yi Wang(王义)61, Yuan Wang(王源)17,31, Z. Wang(王铮)1,58, Z. L.  Wang(王治浪)73, Z. L. Wang(王治浪)2, Z. Q. Wang(王紫祺)12,g, Z. Y. Wang(王至勇)1,64, D. H. Wei(魏代会)14, H. R. Wei43, F. Weidner69, S. P. Wen(文硕频)1, Y. R. Wen(温亚冉)39, U. Wiedner3, G. Wilkinson70, M. Wolke76, C. Wu(吴晨)39, J. F. Wu(吴金飞)1,8, L. H. Wu(伍灵慧)1, L. J. Wu(吴连近)1,64, L. J. Wu(吴连近)19, Lianjie Wu(武廉杰)19, S. G. Wu(吴韶光)1,64, S. M. Wu(吴蜀明)64, X. Wu(吴潇)12,g, X. H. Wu(伍雄浩)34, Y. J. Wu(吴英杰)31, Z. Wu(吴智)1,58, L. Xia(夏磊)72,58, X. M. Xian(咸秀梅)39, B. H. Xiang(向本后)1,64, D. Xiao(肖栋)38,k,l, G. Y. Xiao(肖光延)42, H. Xiao(肖浩)73, Y.  L. Xiao(肖云龙)12,g, Z. J. Xiao(肖振军)41, C. Xie(谢陈)42, K. J. Xie(谢凯吉)1,64, X. H. Xie(谢昕海)46,h, Y. Xie(谢勇 )50, Y. G. Xie(谢宇广)1,58, Y. H. Xie(谢跃红)6, Z. P. Xie(谢智鹏)72,58, T. Y. Xing(邢天宇)1,64, C. F. Xu1,64, C. J. Xu(许创杰)59, G. F. Xu(许国发)1, H. Y. Xu(许皓月)67,2, H. Y. Xu(许皓月)2, M. Xu(徐明)72,58, Q. J. Xu(徐庆君)16, Q. N. Xu30, T. D. Xu(徐腾达)73, W. Xu(许威)1, W. L. Xu(徐万伦)67, X. P. Xu(徐新平)55, Y. Xu(徐月)12,g, Y. Xu(徐月)40, Y. C. Xu(胥英超)78, Z. S. Xu(许昭燊)64, F. Yan(严芳)12,g, H. Y. Yan(闫浩宇)39, L. Yan(严亮)12,g, W. B. Yan(鄢文标)72,58, W. C. Yan(闫文成)81, W. H. Yan(闫文昊)6, W. P. Yan(闫文鹏)19, X. Q. Yan(严薛强)1,64, H. J. Yang(杨海军)51,f, H. L. Yang(杨昊霖)34, H. X. Yang(杨洪勋)1, J. H. Yang(杨君辉)42, R. J. Yang(杨润佳)19, T. Yang(杨涛)1, Y. Yang(杨莹)12,g, Y. F. Yang(杨艳芳)43, Y. H. Yang(杨友华)42, Y. Q. Yang(杨永强)9, Y. X. Yang(杨逸翔)1,64, Y. Z. Yang(杨颖喆)19, M. Ye(叶梅)1,58, M. H. Ye(叶铭汉)8,a, Z. J. Ye(叶子健)56,j, Junhao Yin(殷俊昊)43, Z. Y. You(尤郑昀)59, B. X. Yu(俞伯祥)1,58,64, C. X. Yu(喻纯旭)43, G. Yu13, J. S. Yu(俞洁晟)25,i, L. Q. Yu( 喻丽雯)12,g, M. C. Yu40, T. Yu(于涛)73, X. D. Yu(余旭东)46,h, Y. C. Yu(郁烨淳)81, C. Z. Yuan(苑长征)1,64, H. Yuan(袁昊)1,64, J. Yuan(袁杰)45, J. Yuan(袁菁)34, L. Yuan(袁丽)2, S. C. Yuan(苑思成)1,64, X. Q. Yuan(袁晓庆)1, Y. Yuan(袁野)1,64, Z. Y. Yuan(袁朝阳)59, C. X. Yue(岳崇兴)39, Ying Yue(岳颖)19, A. A. Zafar74, S. H. Zeng63A,63B,63C,63D, X. Zeng(曾鑫)12,g, Y. Zeng(曾云)25,i, Y. J. Zeng(曾溢嘉)1,64, Y. J. Zeng(曾宇杰)59, X. Y. Zhai(翟星晔)34, Y. H. Zhan(詹永华)59, A. Q. Zhang(张安庆)1,64, B. L. Zhang(张伯伦)1,64, B. X. Zhang(张丙新)1, D. H. Zhang(张丹昊)43, G. Y. Zhang(张广义)19, G. Y. Zhang(张耕源)1,64, H. Zhang(张豪)72,58, H. Zhang(张晗)81, H. C. Zhang(张航畅)1,58,64, H. H. Zhang(张宏浩)59, H. Q. Zhang(张华桥)1,58,64, H. R. Zhang(张浩然)72,58, H. Y. Zhang(章红宇)1,58, J. Zhang(张进)81, J. Zhang(张晋)59, J. J. Zhang(张进军)52, J. L. Zhang(张杰磊)20, J. Q. Zhang(张敬庆)41, J. S. Zhang(张家声)12,g, J. W. Zhang(张家文)1,58,64, J. X. Zhang(张景旭)38,k,l, J. Y. Zhang(张建勇)1, J. Z. Zhang(张景芝)1,64, Jianyu Zhang(张剑宇)64, L. M. Zhang(张黎明)61, Lei Zhang(张雷)42, N. Zhang(张楠)81, P. Zhang(张鹏)1,8, Q. Zhang(张强)19, Q. Y. Zhang(张秋岩)34, R. Y. Zhang(张若愚)38,k,l, S. H. Zhang(张水涵)1,64, Shulei Zhang(张书磊)25,i, X. M. Zhang(张晓梅)1, X. Y Zhang40, X. Y. Zhang(张学尧)50, Y.  Zhang(张宇)73, Y. Zhang(张瑶)1, Y.  T. Zhang(张亚腾)81, Y. H. Zhang(张银鸿)1,58, Y. M. Zhang(张悦明)39, Y. P. Zhang(张越鹏)72,58, Z. D. Zhang(张正德)1, Z. H. Zhang(张泽恒)1, Z. L. Zhang(张兆领)34, Z. L. Zhang(张志龙)55, Z. X. Zhang(张泽祥)19, Z. Y. Zhang(张振宇)77, Z. Y. Zhang(张子羽)43, Z. Z.  Zhang(张子扬)45, Zh. Zh. Zhang19, G. Zhao(赵光)1, J. Y. Zhao(赵静宜)1,64, J. Z. Zhao(赵京周)1,58, L. Zhao(赵玲)1, L. Zhao(赵雷)72,58, M. G. Zhao(赵明刚)43, N. Zhao(赵宁)79, R. P. Zhao(赵若平)64, S. J. Zhao(赵书俊)81, Y. B. Zhao(赵豫斌)1,58, Y. L. Zhao(赵艳琳)55, Y. X. Zhao(赵宇翔)31,64, Z. G. Zhao(赵政国)72,58, A. Zhemchugov36,b, B. Zheng(郑波)73, B. M. Zheng(郑变敏)34, J. P. Zheng(郑建平)1,58, W. J. Zheng(郑文静)1,64, X. R. Zheng(郑心如)19, Y. H. Zheng(郑阳恒)64,p, B. Zhong(钟彬)41, C. Zhong(钟翠)19, H. Zhou(周航)35,50,o, J. Q. Zhou(周嘉奇)34, J. Y. Zhou(周佳莹)34, S.  Zhou(周帅)6, X. Zhou(周详)77, X. K. Zhou(周晓康)6, X. R. Zhou(周小蓉)72,58, X. Y. Zhou(周兴玉)39, Y. Z. Zhou(周袆卓)12,g, A. N. Zhu(朱傲男)64, J. Zhu(朱江)43, K. Zhu(朱凯)1, K. J. Zhu(朱科军)1,58,64, K. S. Zhu(朱康帅)12,g, L. Zhu(朱林)34, L. X. Zhu(朱琳萱)64, S. H. Zhu(朱世海)71, T. J. Zhu(朱腾蛟)12,g, W. D. Zhu(朱稳定)41, W. D. Zhu(朱稳定)12,g, W. J. Zhu(朱文静)1, W. Z. Zhu(朱文卓)19, Y. C. Zhu(朱莹春)72,58, Z. A. Zhu(朱自安)1,64, X. Y. Zhuang(庄新宇)43, J. H. Zou(邹佳恒)1, J. Zu(祖健)72,58
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Mongolian Academy of Sciences, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Bristol, H H Wills Physics Laboratory, Tyndall Avenue, Bristol, BS8 1TL, UK
64 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
65 University of Groningen, NL-9747 AA Groningen, The Netherlands
66 University of Hawaii, Honolulu, Hawaii 96822, USA
67 University of Jinan, Jinan 250022, People’s Republic of China
68 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
69 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
70 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
71 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
72 University of Science and Technology of China, Hefei 230026, People’s Republic of China
73 University of South China, Hengyang 421001, People’s Republic of China
74 University of the Punjab, Lahore-54590, Pakistan
75 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
76 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
77 Wuhan University, Wuhan 430072, People’s Republic of China
78 Yantai University, Yantai 264005, People’s Republic of China
79 Yunnan University, Kunming 650500, People’s Republic of China
80 Zhejiang University, Hangzhou 310027, People’s Republic of China
81 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at École Polytechnique Fédérale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Hangzhou 310024, China
1 Introduction
The purely leptonic decays of charmed meson, $D^{+}\to l^{+}\nu_{l}\ (l=e,\ \mu,\ \tau)$, offer an ideal laboratory for testing the Standard Model. These decays offer one of the simplest and most well-understood probes of the $c\to d$ quark flavour-changing transition [1, 2, 3, 4]. However, these decays are subject to helicity suppression, with decay rates proportional to the square of the lepton mass. Consequently, the $D^{+}\to e^{+}\nu_{e}$ decay rate is highly suppressed compared to its muonic [5, 6] and tauonic [5, 7] counterparts, with a branching fraction (BF) expected to be around $10^{-8}$ and beyond the current experimental sensitivity [8, 9].
In the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$, photon emission mitigates helicity suppression, with the strong interaction engaged solely in a single hadronic external state. Theoretically, its differential decay width can be expressed as a function of the photon energy $E_{\gamma}$ in the rest frame of the $D^{+}$ meson [10] via
$\frac{{\rm d}\Gamma}{{\rm d}E_{\gamma}}(D^{+}\to\gamma l^{+}\nu_{l})=$ $\frac{\alpha_{em}G^{2}_{F}\left[F_{V}^{2}(E_{\gamma})+F_{A}^{2}(E_{\gamma})\right]}{6\pi^{2}}$ (1)
×$\times\left|V_{cd}\right|^{2}E_{\gamma}^{3}m_{D^{+}}\left(1-\frac{2E_{\gamma}}{m_{D^{+}}}\right),$
where $\alpha_{em}$ is the fine-structure constant, $G_{F}$ is the Fermi coupling constant, $\left|V_{cd}\right|$ is the Cabibbo-Kobayashi-Maskawa matrix element [11, 12], $m_{D^{+}}$ is the nominal mass of the $D^{+}$ meson [13], and $F_{V}$ and $F_{A}$ are the form factors characterizing the non-perturbative quantum chromodynamics (QCD) dynamics. Various theoretical models have been proposed to investigate the $D^{+}\to\gamma e^{+}\nu_{e}$ decay, including quark models [14, 15, 16], perturbative QCD [17] and QCD factorization approaches [18, 19, 10]. Most theoretical predictions on its BF, $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$, are at the order of $10^{-5}$, as summarized in Table 1.
[…]
The only experimental result of $D^{+}\to\gamma e^{+}\nu_{e}$ comes from the BESIII experiment [20], based on 2.9 $\rm fb^{-1}$ of $e^{+}e^{-}$ annihilation data taken at the center-of-mass energy $\sqrt{s}=3.773\leavevmode\nobreak\ \,\mathrm{GeV}$. The upper limit on the partial BF of $D^{+}\to\gamma e^{+}\nu_{e}$ with $E_{\gamma}>10$ MeV was determined to be $3.0\times 10^{-5}$ at 90% confidence level (C.L.), approaching the range of theoretical predictions [16, 17, 18, 19, 10]. The larger dataset of 20.3 $\rm fb^{-1}$ of $e^{+}e^{-}$ annihilation recently collected at BESIII [21, 22, 23], which includes the previous 2.9  $\rm fb^{-1}$, provides an excellent opportunity to further constrain these theoretical models.
The major challenge in searching for $D^{+}\to\gamma e^{+}\nu_{e}$ is the substantial background contamination from $D^{+}$ semileptonic decays, such as $D^{+}\to\pi^{0}e^{+}\nu_{e}$, $D^{+}\to K^{0}_{L}e^{+}\nu_{e}$ and $D^{+}\to K^{0}_{S}(\to\pi^{0}\pi^{0})e^{+}\nu_{e}$. Photons originating from $\pi^{0}$ decays or long-lived neutral hadrons can deposit energy in the electromagnetic calorimeter, potentially mimicking the radiative photon signal. Conventional cut-based methods have struggled to establish effective discriminators to suppress these backgrounds, whose yields are two orders of magnitude greater than the potential signal.
An unprecedentedly powerful signal identification tool can be accomplished via a deep learning approach, which utilizes a deep neural network (DNN) to distinguish signal from backgrounds. Due to the flexible data representation and modern algorithms, this method exhibits powerful capabilities in recognizing and interpreting underlying relationships and hidden patterns within the topological structure of $e^{+}e^{-}$ annihilation events.
In this work, we present an improved search for $D^{+}\to\gamma e^{+}\nu_{e}$ by analyzing the 20.3 $\mathrm{fb}^{-1}$ data, enhancing signal sensitivity through a deep-learning-based signal identification method. Throughout this paper, charge conjugation is implied by default.
2 BESIII detector and Monte Carlo simulation
The BESIII detector [24] records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [25] in the center-of-mass energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [26, 27, 28]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1\leavevmode\nobreak\ {\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the plastic scintillator TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps, which benefits 85% of the data used in this analysis [29, 30, 31].
Monte Carlo (MC) simulated data samples produced with a geant4-based [32] software package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. A signal MC sample of the decay $D^{+}\to\gamma e^{+}\nu_{e}$ is simulated according to Eq. (1), where the form factors are taken from Ref. [19]. The minimum energy of the radiative photon is set at 10 MeV to avoid the infrared divergence for soft photons. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [33, 34]. The inclusive MC sample includes the production of $D\bar{D}$ pairs (including quantum coherence for the neutral $D$ channels), the non-$D\bar{D}$ decays of the $\psi(3770)$, the ISR production of the $J/\psi$ and $\psi(3686)$ states, and the continuum processes incorporated in kkmc [33, 34]. All particle decays are modelled with evtgen [35, 36] using branching fractions either taken from the Particle Data Group [13], when available, or otherwise estimated with lundcharm [37, 38]. Final state radiation from charged final state particles is incorporated using the photos package [39].
3 Initial data analysis
Taking advantage of the $D^{+}D^{-}$ pair production in $e^{+}e^{-}$ annihilation at $\sqrt{s}=3.773$ GeV, we employ a double-tag (DT) analysis method [40] to constrain the kinematics of $D^{+}\to\gamma e^{+}\nu_{e}$ and suppress non-$D^{+}D^{-}$ backgrounds. The $D^{-}$ meson is firstly reconstructed through six specific hadronic decay modes, $K^{+}\pi^{-}\pi^{-}$, $K^{+}\pi^{-}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{-}$, $K^{0}_{S}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{+}\pi^{-}\pi^{-}$ and $K^{+}K^{-}\pi^{-}$. Here, $K^{0}_{S}$ is reconstructed with a $\pi^{+}\pi^{-}$ pair and $\pi^{0}$ is reconstructed from two photons. These decay modes are referred to as tag modes, and the selected events are called single-tag (ST) events. In the presence of ST $D^{-}$ meson, the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ is then searched for in its recoil system, with selected events referred to as DT events. The BF of the signal decay is determined by
$\mathcal{B}_{\rm sig}=\frac{N_{\rm DT}}{\sum_{i}\left(N_{\rm ST}^{i}\epsilon_{\rm DT}^{i}/\epsilon^{i}_{\rm ST}\right)},$ (2)
where $N_{\rm DT}$ is sum of DT yields for all tag modes, and $N^{i}_{\rm ST}$, $\epsilon^{i}_{\rm ST}$, and $\epsilon^{i}_{\rm DT}$ are the ST yield, ST efficiency, and DT efficiency for the tag mode $i$, respectively.
[…]
The $K_{S}^{0}$ candidates are reconstructed from two oppositely charged tracks satisfying $V_{z}<$ 20 cm. The two charged tracks are assigned as $\pi^{+}\pi^{-}$ without imposing further PID criteria. They are constrained to originate from a common vertex and are required to have an invariant mass in the range of (487, 511) $\,\mathrm{MeV}/c^{2}$ corresponding to approximately ±$\pm 3\sigma$ of the detection resolution. The decay length of the $K^{0}_{S}$ candidate is required to be greater than twice the vertex resolution away from the IP.
Photon candidates are identified using showers in the EMC. The deposited energy of each shower must be more than 25 MeV in the barrel region ($|\cos\theta|<0.80$) and more than 50 MeV in the end cap region ($0.920.86<|\cos\theta|<$ 0.92). To exclude showers that originate from charged tracks, the angle subtended by the EMC shower and the position of the closest charged track at the EMC must be greater than 10 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns. The $\pi^{0}$ mesons are reconstructed using a kinematic fit from all combinations of two photon candidates by constraining their invariant mass to the nominal $\pi^{0}$ mass [13]. The invariant mass of two photons before the kinematic fit is required to be in the range of (115, 150) $\,\mathrm{MeV}/c^{2}$. The combinations with both photons from end cap EMC regions are rejected due to worse resolution.
After selecting the component particles, we reconstruct ST $D^{-}$ meson using their combinations. Two kinematic variables, the energy difference $\Delta E$ and the beam-constrained mass $M_{\rm BC}$, are calculated for signal extraction
$\Delta E$ $=E_{D^{-}}-E_{\rm beam},$ (3)
$M_{\rm BC}$ $=\sqrt{E^{2}_{\rm beam}/c^{4}-\left|\vec{p}_{D^{-}}\right|^{2}/c^{2}},$ (4)
where $E_{\rm beam}$ is the beam energy, $\vec{p}_{D^{-}}$ and $E_{D^{-}}$ are the momentum vector and energy of $D^{-}$ candidate in the $e^{+}e^{-}$ rest frame, respectively. If multiple combinations are present in an event, the combination with the smallest $|\Delta E|$ is retained for further analysis for each tag mode and charge. To suppress combinatorial backgrounds, mode-dependent requirements on the $\Delta E$ are imposed on the accepted candidates within approximate ±$\pm 3.5\sigma$ range of the detection resolution, as listed in Table 3.
[…]
The $\Delta E$ requirements, ST yields in data, ST and DT efficiencies for each tag mode in data. Note the efficiencies do not include the BFs of $\pi^{0}$ and $K^{0}_{S}$ subsequent decays. Tag mode $\Delta E$ (MeV) $N_{\rm{ST}}^{i}\ (\times 10^{3})$ $\epsilon_{\rm{ST}}^{i}\ (\%)$ $\epsilon_{\rm{DT}}^{i}\ (\%)$ $K^{+}\pi^{-}\pi^{-}$ (25, 24)2524(-25,\ 24)( - 25 , 24 ) $5552.8\pm 2.5$ $51.10$ 10.78 $K^{0}_{S}\pi^{-}$ (25, 26)2526(-25,\ 26)( - 25 , 26 ) $656.5\pm 0.8$ $51.42$ 10.69 $K^{+}\pi^{-}\pi^{-}\pi^{0}$ (57, 46)5746(-57,\ 46)( - 57 , 46 ) $1723.7\pm 1.8$ $24.40$ 5.42 $K^{0}_{S}\pi^{-}\pi^{0}$ (62, 49)6249(-62,\ 49)( - 62 , 49 ) $1442.4\pm 1.5$ $26.45$ 5.57 $K^{0}_{S}\pi^{-}\pi^{-}\pi^{+}$ (28, 27)2827(-28,\ 27)( - 28 , 27 ) $790.2\pm 1.1$ $29.59$ 6.62 $K^{+}K^{-}\pi^{-}$ (24, 23)2423(-24,\ 23)( - 24 , 23 ) $481.4\pm 0.9$ $40.91$ 8.96
[…]
Fits to the $M_{\rm BC}$ distributions of ST $D^{-}$ candidates for each tag mode in data. The black dots with error bars represent data. The blue curves are the fit results. The red dotted curves are the fitted combinatorial background shapes. The yellow histograms are the MC simulated background scaled to the luminosity of data. The red arrows indicate signal region.
The signal candidates for $D^{+}\to\gamma e^{+}\nu_{e}$ are selected from the remaining charged tracks and showers, excluding those associated with the ST $D^{-}$ meson. We require only one track left with charge opposite to that of the ST $D^{-}$ meson, following the same selection criteria as those in the tag side. This track must be identified as a positron combining the PID information from $\mathrm{d}E/\mathrm{d}x$, TOF and EMC, satisfying $\mathcal{L}(e)>0.001$ and $\mathcal{L}(e)/[\mathcal{L}(e)+\mathcal{L}(\pi)+\mathcal{L}(K)]>0.8$ for the combined confidence levels. We also require at least one photon candidate satisfying the same selection criteria as used in the tag side. If multiple photon candidates pass the selection, the most energetic one is chosen as the signal radiative photon. To improve the kinematic resolution of the positron, final-state-radiation recovery is performed by adding the momenta of any remaining photon candidates back to the reconstructed momentum of positron candidate if their opening angles with respect to the IP are less than $5^{\circ}$.
To derive the information of the missing neutrino, a kinematic observable $U_{\rm miss}$ is defined as
$U_{\rm miss}=E_{\rm miss}-\left|\vec{p}_{\rm miss}\right|c,$ (5)
[…]
We also study $D^{+}\to\pi^{0}e^{+}\nu_{e}$ as a validation channel in this work. In reconstructing the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ events, additional DT selection criteria are applied after those mentioned above. At least one $\pi^{0}$ candidate is required, in which the one with minimum $\chi^{2}$ of the kinematic fit is chosen as the signal $\pi^{0}$ candidate. Another kinematic observable, $U^{\prime}_{\rm miss}$, is defined as
$U^{\prime}_{\rm miss}=E^{\prime}_{\rm miss}-\left|\vec{p^{\prime}}_{\rm miss}\right|c,$ (6)
where $E^{\prime}_{\rm miss}=E_{\rm beam}-E_{e^{+}}-E_{\pi^{0}}$ and $\vec{p^{\prime}}_{\rm miss}=\vec{p}_{D^{+}}-\vec{p}_{e^{+}}-\vec{p}_{\pi^{0}}$ are calculated using the energy and momentum of the signal $\pi^{0}$ candidate, respectively.
Figure 3(a) presents the $U_{\rm miss}$ distributions for both data and inclusive MC samples after applying the above selections. The major backgrounds are from $D^{+}\to\pi^{0}e^{+}\nu_{e}$, $D^{+}\to K^{0}_{L}e^{+}\nu_{e}$, and $D^{+}\to K^{0}_{S}(\to\pi^{0}\pi^{0})e^{+}\nu_{e}$. Notably, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background exhibits a non-trivial shape within the signal region. The high background level makes it difficult to re-
[…]
(a) The $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) the $U^{\prime}_{\rm miss}$ distribution of the validation channel $D^{+}\to\pi^{0}e^{+}\nu_{e}$ after the DT selection. The black dots with error bars represent data. The histograms with filled colors represent different background components in the inclusive MC sample, respectively. The red hollow histogram indicates an arbitrarily scaled signal shape.
liably extract the signal yield from the $U_{\rm miss}$ distribution with adequate sensitivity, thereby requiring further signal identification treatments as elaborated below.
4 Deep-learning-based signal identification
The main challenge in distinguishing the signals from massive backgrounds lies in extracting distinctive features from the data. Notably, the shower patterns observed in the EMC provide key information for this task. Background events involving a $\pi^{0}$ typically result in a higher number of photon candidates than the signal, whereas $K^{0}_{L}$-induced showers often exhibit more dispersed cluster shapes compared to those from photons [42]. Additionally, the kinematics of the particles, influenced by the unique phase spaces and dynamics of their decays, offer another aspect of differentiation. Conventional cut-based methods characterize such distinctions via hand-crafted variables derived from practical understanding of the involved physics, such as the number of $\pi^{0}$-like photon combinations, the lateral moment of photon candidates [20], and the $\chi^{2}$ values from kinematic fits under specific final-state hypotheses [43]. In contrast, deep learning methods aim to process minimally refined data at the level of fundamental detector responses to the penetrations of the final state particles. The DNN is expected to automatically explore underlying correlations and recognize distinctive information beyond prior knowledge.
The dataset for training the DNN is constructed by randomly shuffling signal and background events sourced from the relevant MC samples. The events are categorized into three groups with equal statistics: the $D^{+}\to\gamma e^{+}\nu_{e}$ signal, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background, and other inclusive backgrounds. The total training dataset contains approximate 3.3 million events, with 80% used for training and the remaining 20% reserved for validation. The input information from an event to DNN includes all charged tracks reconstructed in the MDC and isolated showers clustered in the EMC. These particles are organized as a point cloud [44] structure, i.e., an unordered and variable-sized set of the points in a high-dimensional feature space. For each charged track, the features comprise the azimuthal and polar angles in the center-of-mass frame, the charge, the magnitude of momentum, and the parameters characterizing its helical trajectory in the MDC. Additionally, low-level measurements from the MDC, TOF, and EMC are incorporated as implicit information for particle identification and reconstruction quality. For each shower, the features consist of the azimuthal and polar angles, the energy deposition, the count of fired crystals, the time measurement, and the parameters describing its expansion scope among nearby crystals. A full list of input features is provided in Table 4, and these features are found to be consistent between data and MC simulation.
[…]
The input particle feature variables and their availability in tracks and showers. If any measurement is unavailable from the detector readouts, the default value is set at zero. Variable Definition Track Shower $\theta$ azimuthal angle $\phi$ polar angle $p$ momentum magnitude charge electric charge $d_{\rho}$ distance in the $x-y$ plane from the point of closest approach (POCA) of the track to the IP $d_{Z}$ distance along $z$-axis from the POCA of the track to the IP $\phi_{0}$ azimuthal angle of the track at the POCA $\kappa$ scaled inverse of the track’s momentum $\tan\lambda$ tangent value of the track’s angle at the POCA with respect to the $x-y$ plane $\chi_{{\rm d}E/{\rm d}x}(e)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated positron value $\chi_{{\rm d}E/{\rm d}x}(\pi)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated pion value $\chi_{{\rm d}E/{\rm d}x}(K)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated kaon value $\chi_{\rm TOF}(\pi)$ deviation in TOF measurement of the track from the anticipated pion value $\chi_{\rm TOF}(K)$ deviation in TOF measurement of the track from the anticipated kaon value $Q_{\rm TOF}$ pulse height of the track in TOF corresponding to its deposited energy $\beta_{\rm TOF}$ velocity of the track measured by TOF $E_{\rm raw}$ total deposited energy in EMC $E_{\rm seed}$ deposited energy of the crystal at the shower center $E_{3\times 3}$ sum of energies of $3\times 3$ crystal matrices surrounding the shower center $E_{5\times 5}$ sum of energies of $5\times 5$ crystal matrices surrounding the shower center $N_{\rm hits}$ count of fired crystals in the shower time EMC time measurement secondary moment shower expansion parameter defined as $\sum_{i}E_{i}r^{2}_{i}/\sum_{i}E_{i}$ lateral moment shower expansion parameter defined as $\sum_{i=3}^{n}E_{i}r_{i}^{2}/(E_{1}r_{0}^{2}+E_{2}r_{0}^{2}+\sum_{i=3}^{n}E_{i}r_{i}^{2})$ $A_{20}$ moment shower expansion parameter defined as $\sum_{i}\frac{E_{i}}{E_{tot}}f_{2,0}(\frac{r_{i}}{R_{0}})$, $f_{2,0}(x)=2x^{2}-1$ $A_{42}$ moment same but defined as $\left|\sum_{i}\frac{E_{i}}{E_{tot}}f_{4,2}(\frac{r_{i}}{R_{0}})e^{2i\phi}\right|$$f_{4,2}(x)=4x^{4}-3x^{2}$
The architecture of the DNN largely inherits from the Particle Transformer (ParT) [45], with adaptations tailored for the BESIII experiment. A key innovation of the ParT is the injection of particle physics knowledge into the Transformer network [46], where a set of pairwise interaction features are calculated using the energy-momentum vectors for each pair of particles and incorporated as biases $\mathbf{U}$ in the self-attention mechanism. This treatment efficiently captures the kinematic relationships between particles, significantly improving the ParT performance. For the BESIII application, we account for its unique properties, such as the purely spherical coordinate system and energy-momentum conservation in 3D space, by designing a new set of pairwise interaction features $\mathbf{U}_{\rm BES}$ as
$\mathbf{U}_{\rm BES}=\left\{\begin{aligned} &1-\cos\theta_{a,b},\\ &{\rm min}\left(E_{a},\,E_{b}\right)\left(1-\cos\theta_{a,b}\right),\\ &{\rm min}\left(E_{a},\,E_{b}\right)/\left(E_{a}+E_{b}\right),\\ &\left(E_{a}+E_{b}\right)^{2}-\lVert\vec{p}_{a}+\vec{p}_{b}\rVert^{2},\end{aligned}\right.$ (7)
where $E_{a}$ and $E_{b}$ are the energies of the two particles, $\vec{p}_{a}$ and $\vec{p}_{b}$ are their momentum vectors, and $\theta_{a,b}$ is the opening angle between these vectors in the center-of-mass frame. For the model hyperparameters, this adaptation includes four particle attention blocks and two class attention blocks. It uses a particle embedding encoded from the input particle features via a 3-layer Multilayer Perceptron with 128, 256, 128128256128128,\ 256,\ 128128 , 256 , 128 nodes in each layer, and an interaction embedding encoded from the input pairwise features using a 4-layer pointwise 1D convolution with 64, 64, 64, 8646464864,\ 64,\ 64,\ 864 , 64 , 64 , 8 channels. A dropout rate of 0.1 is applied to prevent overfitting.
The training of the DNN is performed on the aforementioned dataset for 50 epochs with a batch size of 512 and an initial learning rate of 0.001. An additional treatment is employed during training to address the potential correlation between the DNN output and the $U_{\rm miss}$ distribution, which could inadvertently create signal-like structures in the $U_{\rm miss}$ distribution of remaining background events after applying the selection criteria on the DNN output [47, 48, 49, 50]. To decouple the DNN output from the $U_{\rm miss}$ observable, we employ an iterative approach by assigning a weight $\omega(U_{\rm miss})$ to each background event which adjusts its contribution to the loss function. Events with higher weights have a greater impact on the loss function, thus being better classified during training. The weight $\omega_{i}^{j}(U_{\rm miss}^{j})$ for the event $j$ in the $i$-th iteration is calculated as
$\omega_{i}^{j}(U_{\rm miss}^{j})=\omega_{i-1}^{j}(U_{\rm miss}^{j})\left[\left.\frac{p^{\rm BKG}_{i-1}(U_{\rm miss})}{p^{\rm BKG}_{\rm orig}(U_{\rm miss})}\right|_{U_{\rm miss}=U_{\rm miss}^{j}}\right],$ (8)
where $\omega_{0}^{j}(U_{\rm miss}^{j})=1$, $p^{\rm BKG}_{i-1}(U_{\rm miss})$ represents the normalized probability density function for the remaining background shape after the $(i-1)$-th iteration, and $p^{\rm BKG}_{\rm orig}(U_{\rm miss})$ denotes the original background shape before applying the DNN output veto. For each iteration, the DNN is re-trained, and both $p^{\rm BKG}_{i-1}(U_{\rm miss})$ and $\omega_{i}^{j}(U_{\rm miss}^{j})$ are updated. Intuitively, the process converges when $p^{\rm BKG}_{i-1}(U_{\rm miss})$ approaches $p^{\rm BKG}_{\rm orig}(U_{\rm miss})$, resulting in a trivial background shape that facilitates signal extraction. In practice as illustrated in Figure 4, this iterative approach partially corrects the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background shape, while successfully retrieves the original for other inclusive background shape. This result demonstrates both its effectiveness and limitations in addressing the mass correlation issue. Furthermore, another machine-learning technique called model ensemble is employed. With some randomness factors incorporated in the training such as network initialization, batch processing sequence and dropout [51], we train a total of 20 DNNs in parallel in each iteration. The outputs of these DNNs are averaged for each event during inference, creating an ensemble DNN that offers better robustness and generalization than any single DNN model.
[…]
Normalized (a) $D^{+}\to\pi^{0}e^{+}\nu_{e}$ and (b) other inclusive background shapes in the $U_{\rm miss}$ distribution during the iteration process.
The DNN output assigns three scores between [0, 1]01[0,\,1][ 0 , 1 ] to each event, reflecting the probabilities of the event belonging to the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other background categories. Since the sum of the three scores always equals one, only two of them are independent. As the final step in the event selection, we require the score for the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background to be less than 0.15 and the score for other backgrounds to be less than 0.05. These cut values are optimized using the Punzi significance metric, $\epsilon/(1.5+\sqrt{B})$ [52], where $\epsilon$ represents the signal efficiency and $B$ is the expected background yield. Figure 4(a) illustrates the $U_{\rm miss}$ distribution after implementing the deep learning approach. The backgrounds are reduced by more than two orders of magnitude at the cost of approximately two-thirds of the signal efficiency, significantly enhancing the sensitivity of the signal search.
[…]
(a) The $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) the $U^{\prime}_{\rm miss}$ distribution of the validation decay $D^{+}\to\pi^{0}e^{+}\nu_{e}$ after the DNN vetoes. The black dots with error bars represent data. The histograms with filled colors represent different background components in the inclusive MC sample, respectively. The red hollow histogram indicates an arbitrarily scaled signal shape.
5 Signal extraction
According to Eq. (2), the DT efficiencies for each tag mode, $\epsilon^{i}_{\rm DT}$, are estimated using signal MC sample and are summarized in Table 3. To obtain the DT yield $N_{\rm DT}$ in data, a maximum likelihood fit is performed on the $U_{\rm miss}$ distribution combining all the tag modes. The fit includes three components, namely, the signal, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other backgrounds. The shapes for these components are taken from the corresponding MC simulations, and their yields are all floated in the fit. The fit gives $N_{\rm DT}=12.2\pm 8.6$ and $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})=(0.54\pm 0.38)\times 10^{-5}$ as shown in Figure 5(a), where the uncertainties are statistical only. Since no significant signal is observed, an upper limit on the signal BF will be set after accounting for all systematic uncertainties.
[…]
Fits to (a) the $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) the $U^{\prime}_{\rm miss}$ distribution of the validation decay $D^{+}\to\pi^{0}e^{+}\nu_{e}$. The points with error bars represent data. The blue line represents the fitted curve. The red dashed area represents signal component and other dashed lines represent background components.
6 Validation
A common concern in adopting the deep learning method in particle physics experiments is the coupling between potential analysis biases and the sophisticated DNN structure, which makes these biases difficult to validate, quantify and calibrate. The extensive and diverse datasets collected at BESIII facilitate various control samples in data to investigate these issues. In this study, our primary background process $D^{+}\to\pi^{0}e^{+}\nu_{e}$ serves as a natural control sample. By partially reversing the DNN vetoes, i.e., to require the score for the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background to be greater than 0.15 while keeping the score for other backgrounds below 0.05, we can isolate a $D^{+}\to\pi^{0}e^{+}\nu_{e}$ sample with a purity of over 98% and measure its absolute BF. Figures 3(b) and 4(b) illustrate the $U^{\prime}_{\rm miss}$ distributions before and after applying the partially reversed DNN vetoes, with significant subtraction of backgrounds other than $D^{+}\to\pi^{0}e^{+}\nu_{e}$.
To measure the absolute BF of $D^{+}\to\pi^{0}e^{+}\nu_{e}$, we use the $\epsilon^{i}_{\rm DT}$ estimated from MC simulations and extract the $N_{\rm DT}$ via a maximum likelihood fit shown in Figure 5(b). In the fit, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ signal is modeled using the MC simulated shape convolved with a Gaussian resolution function. The other background component is described with the inclusive MC shape with the yield floated. The final fit to all data gives $N_{\rm DT}=3004\pm 59$ and $\mathcal{B}(D^{+}\to\pi^{0}e^{+}\nu_{e})=(3.629\leavevmode\nobreak\ \pm 0.071)\leavevmode\nobreak\ \times 10^{-3}$ where the uncertainty is statistical only. This result is consistent with the previous BESIII measurement $(3.63\leavevmode\nobreak\ \pm 0.08_{\rm stat.}\leavevmode\nobreak\ \pm 0.05_{\rm syst.})\times 10^{-3}$ [53] based on about one sevenths of full data, thereby validating the reliability of our DNN method. The residual deviations are considered as a source of systematic uncertainty, as detailed in the next section.
7 Systematic uncertainties
The systematic uncertainties in the BF determination stem from various sources, with many related to the ST selection being cancelled benefiting from the DT technique. The remaining systematic sources can be grouped into two categories: multiplicative uncertainties, which affect the efficiency, and additive uncertainties, which impact the signal yield. These two types of uncertainties contribute to the upper limit of the BF through different approaches.
The multiplicative uncertainty sources include the ST yield, positron tracking, positron PID, photon reconstruction, MC model, and DNN vetoes, as summarized in Table 7. The uncertainty due to ST yield is estimated by varying the descriptions of signal and background components in the $M_{\rm BC}$ fits and is assigned as 0.3%. A control sample of radiative Bhabha scattering events is used to study the position tracking and PID uncertainties, and their values are assigned as 1.0% for both [54]. The photon reconstruction uncertainty is assigned to be 1.0% using a control sample of $J/\psi\to\pi^{+}\pi^{-}\pi^{0},\,\pi^{0}\to\gamma\gamma$ [55]. The uncertainty related to MC modeling is assigned as the difference in the signal efficiencies of the nominal MC model with an alternative single-pole model [18], which is 9.1%.
Special considerations are given to systematic uncertainties related to the DNN vetoes [56]. As suggested by some frontier surveys [13, 57], we consider two primary uncertainty sources: model uncertainty and domain shift. Model uncertainty arises from our limited knowledge of the best model and is quantified by the relative shift in signal efficiency across the 20 DNNs trained during the model ensemble process. The standard deviation of these efficiencies, divided by their average, 5.3%, is taken as the corresponding systematic uncertainty. Domain shift describes the mismatch between datasets for training and inference, reflecting potential data-MC inconsistencies. This uncertainty is evaluated using the control sample of $D^{+}\to\pi^{0}e^{+}\nu_{e}$ described in Section 6, with the deviation between our BF measurement and the previous BESIII measurement [53] calculated to be 2.2%, combining the center value shift and $1\sigma$ statistical uncertainty. In total, the multiplicative systematic uncertainty is determined to be 10.9% by adding the above sources in quadrature.
[…]
The multiplicative systematic uncertainties. Source Size (%) ST yield 0.3 Positron tracking 1.0 Positron PID 1.0 Photon reconstruction 1.0 MC model 9.1 DNN vetoes 5.8 Total 10.9
The additive systematic uncertainties arise from the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other background shapes used in the DT fit. These effects are evaluated by testing alternative descriptions of the shapes. One variation is using parameterized functions, where the signal and $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background shapes are modeled by double-Gaussian functions, and the other background shape by a 3rd-order Chebyshev polynomial. Another change is using the MC-simulated shapes before the final iteration of the mass decorrelation process, as well as across the 20 DNN predictions in the model ensemble process.
8 Results
As no significant signal is observed in data, the upper limit on $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})$ is set using a Bayesian method described in Ref. [58]. We perform a series of maximum likelihood fits as described in Section 5, with the signal yield $N_{\rm DT}$ fixed to each scanned value. The corresponding maximum likelihood values are used to construct a discrete likelihood distribution, $\mathcal{L}(\mathcal{B}_{\rm sig})$. The systematic uncertainties are incorporated in two steps: first, the additive uncertainties are addressed by varying the DT fit method, with the most conservative upper limit retained. Next, the multiplicative uncertainties are accounted for by smearing the likelihood distribution as [59]
$\mathcal{L}^{\prime}(\mathcal{B}_{\rm sig})\propto\int^{1}_{0}\mathcal{L}\Big{(}\mathcal{B}_{\rm sig}\cdot\frac{\epsilon}{\epsilon_{0}}\Big{)}e^{-\frac{(\epsilon-\epsilon_{0})^{2}}{2\sigma_{\epsilon}^{2}}}\mathrm{d}\epsilon,$ (9)
where $\epsilon_{0}$ is the nominal signal efficiency and $\sigma_{\epsilon}$ is the multiplicative uncertainty corresponding to the efficiency value. By integrating the $\mathcal{L}^{\prime}(\mathcal{B}_{\rm sig})$ curve up to 90% of the area for $N_{\rm DT}>0$ and calculating the corresponding BF using Eq. (2), we set the upper limit on the BF of $D^{+}\to\gamma e^{+}\nu_{e}$ at 90% C.L. to be $1.2\times 10^{-5}$, as shown in Figure 8,
[…]
The normalized likelihood distributions before and after incorporating the multiplicative and additive systematic uncertainties. The blue dashed line represents the initial distribution and the red solid line represents the distribution with systematic uncertainties considered. The red arrow indicates the final upper limit.
9 Summary
In conclusion, we have conducted an improved search for the radiative leptonic decay $D^{+}\to\gamma e^{+}\nu_{e}$ using 20.3 $\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at $\sqrt{s}=3.773$ GeV with the BESIII detector. Our investigation does not reveal a significant signal, and establish an upper limit on its BF with radiative photon energy over 10 MeV to be $1.2\times 10^{-5}$ at 90% C.L.. This result excludes most existing theoretical predictions [16, 17, 18, 19, 10], thereby indicating the necessity of further investigation of the decay mechanism. The key advancement of this study is the application of a novel deep learning approach, which effectively distinguishes the signals from massive backgrounds. Our results showcase the potential and wide-ranging applicability of deep learning techniques in particle physics research.